A bond investor is considering two 10 year maturity bonds both rated AA:  
the municipal bond is yielding 2.47% and the corporate bond is yielding 4.36%.  
At what marginal tax rate would the bond investor be indifferent between the two bonds?
Answer #1

Aftertax corporate bond yield=4.36*(1-tax rate)

Hence for investor to be indifferent:

4.36*(1-tax rate)=2.47%

1-tax rate=2.47/4.36

tax rate=1-(2.47/4.36)

which is equal to

=43.35%(Approx).